Sally is creating a rectangular quilt and the length is equal to five inches less than the width. The perimeter is 30 inches. What are the dimensions of the quilt?

Answer by Alan3354(69443)

P = 2L + 2W = 30
L + W = 15
L = W - 5
-----
W-5 + W = 15
2W = 20
W = 10
L = 5
----------------
That's a small quilt, big enough for a rat maybe.